OpenStudy (anonymous):

On the graph of the equation 5x – 2y = 20, what is the value of the y-intercept?

OpenStudy (saifoo.khan):

When y-intercepts, x =0 5(0) - 2y = 20

OpenStudy (anonymous):

5x+2y=20 is the same as y=10-2.5x Put values in for x and plot eg. x=0, y = 10 x=1, y = 7.5 x=2, y = 5 ad naseum

OpenStudy (anonymous):

The options are 4, -4, 10 or -10... And i'm not getting either of those..

OpenStudy (anonymous):

Just use Saifoo's explanation and solve for "y" because at the y-intercept, x equals zero because that is where the line intersects the y axis.

OpenStudy (anonymous):

I did, but it still didn't work...

OpenStudy (anonymous):

5(0) - 2y = 20 0-2y = 20 -2y=20 solve for y

OpenStudy (anonymous):

Okay, I got it now... thank you :)